OpenStudy (anonymous):

using euclid's division algorithm,find the HCF of 954 and 1638.

OpenStudy (anonymous):

anyone plz answer ma question.

OpenStudy (maheshmeghwal9):

ya I am calculating.wait please

OpenStudy (anonymous):

k

OpenStudy (maheshmeghwal9):

you know that euclid's division lemma is dividend=divisor*quotient+remainder

OpenStudy (maheshmeghwal9):

so,1638=954*1+684

OpenStudy (maheshmeghwal9):

since ,remainder is not zero.therefore we take divisor as dividend and remainder as divisor

OpenStudy (maheshmeghwal9):

now, 954=684*1+270

OpenStudy (maheshmeghwal9):

since remainder is non-zero therefore we follow that step again now, 684=270*2+144

OpenStudy (maheshmeghwal9):

again, 270=144*1+126

OpenStudy (maheshmeghwal9):

again, 144=126*1+18

OpenStudy (maheshmeghwal9):

again & last step; 126=18*7+0.

OpenStudy (anonymous):

so at last the remainder will be zero?

OpenStudy (maheshmeghwal9):

Since the remainder is zero therefore we got the H.C.F. of 954 & 1638; i.e.,7.

OpenStudy (maheshmeghwal9):

did you understand it?

OpenStudy (anonymous):

yss bt d 3rd step after dt y u r calculating again? i mean 270*2+144

OpenStudy (maheshmeghwal9):

I told you that remainder was not zero.That's why.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

can a quadratic polynomial have no zero?

OpenStudy (maheshmeghwal9):

More problems, than ask me.

OpenStudy (anonymous):

k

OpenStudy (maheshmeghwal9):

what are you asking. I don't understand.

OpenStudy (maheshmeghwal9):

about quadratic polynomial?

OpenStudy (anonymous):

yss

OpenStudy (maheshmeghwal9):

ya a quadratic polynomial has zeroes as its domain but when quadratic polynomial is equated to zero,It has zeroes as its roots

OpenStudy (anonymous):

kk

OpenStudy (maheshmeghwal9):

did u understand it?very intelligent!

OpenStudy (anonymous):

say yss or no 4 d question can a quadratic polynomial hav no zero?

OpenStudy (anonymous):

can uuuuuuuuuuuuuuu

OpenStudy (maheshmeghwal9):

ya

OpenStudy (anonymous):

k